OpenStudy (anonymous):

Are the lines y = -x - 2 and 4x + 4y = 16 perpendicular? Explain. Choose one answer. a. Yes; their slopes are equal. b. Yes; their slopes have product -1. c. No; their slopes are not equal d. No; their slopes are not opposite reciprocals.

OpenStudy (littlebird):

no

OpenStudy (littlebird):

d

OpenStudy (anonymous):

i have 4 others can you please help me with them!

OpenStudy (littlebird):

maybe

OpenStudy (perl):

Please guide the student, don't give the answer.

OpenStudy (littlebird):

sure

OpenStudy (anonymous):

Is the line through points P(-3, -2) and Q(2, 3) perpendicular to the line through points R(10, -1) and S(15, -6)? Explain. Choose one answer. a. No; their slopes are not equal. b. No, their slopes are not opposite reciprocals. c. Yes; their slopes are equal. d. Yes; their slopes have product -1.

OpenStudy (littlebird):

y=mx+b m is the slope You have to use the formula (y-y)/(x-x) to find the slope if one of the slopes of the two equations is a negative reciprocal, then they are perpendicular

OpenStudy (littlebird):

Perpendicular lines are a bit more complicated. If you visualize a line with positive slope (so it's an increasing line), then the perpendicular line must have negative slope (because it will be a decreasing line). So perpendicular slopes have opposite signs. The other "opposite" thing with perpendicular slopes is that their values are reciprocals; that is, you take the one slope value, and flip it upside down. Put this together with the sign change, and you get that the slope of the perpendicular line is the "negative reciprocal" of the slope of the original line --- and two lines with slopes that are negative reciprocals of each other are perpendicular to each other. In numbers, if the one line's slope is m = 4/5, then the perpendicular line's slope will be m = -5/4. If the one line's slope is m = -2, then the perpendicular line's slope will be m = 1/2.

OpenStudy (anonymous):

That was a bit confusing. I'm not the best in math.

OpenStudy (anonymous):

I kinda want to say a, but I have a feeling i'm absolutely wrong.

OpenStudy (littlebird):